NON STATISTICAL HYPERFINE POPULATION DISTRIBUTION OF I(^{2}P_{1/2}) UPON PHOTOLYSIS OF I_{2} VIA THE B^{3}Π(O_{u}^{+}) STATE

\underline{T. VAN MARTER} and M. C. HEAVEN, Department of Chemistry, Emory University, Altlanta, GA 30322.

A non statistical hyperfine population distribution was detected for I(^{2}P_{1/2}) after photolysis of I_{2} via the B^{3}Π(O_{u}^{+}) state. Dissociation of a homonuclear diatomic from an unperturbed electronic state will result in a statistical population of the atomic hyperfine levels. For direct dissociation of I_{2}(B), the I(^{2}P_{1/2}) F=2 and F=3 levels are expected to show a population ratio of 5/7 (N(F=2)/N(F=3)). However, an investigation of the hyperfine population distribution on the photolysis energy revealed a strong dependence with population ratios reaching 8/1 with photolysis energies near the I_{2}(B) dissociation limit. It is believed that avoided crossings between the potential energy curves of I_{2} at large internuclear distances are responsible for the non statistical hyperfine population distribution.

Work supported by the Air Force Office of Scientific Research.